Documenta Mathematica

Extra Volume: Alexander S. Merkurjev’s Sixtieth Birthday, 2015

Preface  1
Merkurjev’s Faves  5

Aravind Asok and Jean Fasel Secondary Characteristic Classes

and the Euler Class  7–29

Asher Auel, R. Parimala, and V. Suresh

Quadric Surface Bundles over Surfaces  31–70

E. Bayer-Fluckiger, V. Emery, and J. Houriet Hermitian Lattices and Bounds in K-Theory

of Algebraic Integers  71–83

Mikhail Borovoi and Boris Kunyavskiĭ

Stably Cayley Semisimple Groups  85–112

Baptiste Calmès, Kirill Zainoulline, Changlong Zhong

Equivariant Oriented Cohomology of Flag Varieties  113–144

Denis-Charles Cisinski, Frédéric Déglise

Integral Mixed Motives in Equal Characteristic  145–194

Jean-Louis Colliot-Thélène Descente galoisienne sur le second groupe de Chow :

mise au point et applications  195–220

Haruzo Hida

Limit Mordell–Weil Groups and their p-Adic Closure 221–264

Detlev W. Hoffmann Motivic Equivalence

and Similarity of Quadratic Forms  265–275

Bruno Kahn, R. Sujatha Birational Geometry and Localisation of Categories With Appendices by Jean-Louis Colliot-Thélène

and Ofer Gabber  277–334

D. Kaledin

K-Theory as an Eilenberg-Mac Lane Spectrum  335–365

Nikita A. Karpenko Minimal Canonical Dimensions

of Quadratic Forms  367–385

Max-Albert Knus and Jean-Pierre Tignol

Triality and algebraic groups of type ^{3}D_{4}  387–405

Marc Levine, Girja Shanker Tripathi Quotients of MGL,

Their Slices and Their Geometric Parts  407–442

Roland Lötscher Essential Dimension of Separable Algebras

Embedding in a Fixed Central Simple Algebra  443–459

A. S. Merkurjev Divisible Abelian Groups are Brauer Groups (Translation of an article originally published in Russian in

Uspekhi Mat. Nauk, vol. 40 (1985), no. 2(242), 213–214)  461–463

Matthew Morrow Zero Cycles on Singular Varieties

and Their Desingularisations  465–486

Manuel Ojanguren

Wedderburn’s Theorem for Regular Local Rings  487–490

I. Panin, V. Petrov Rationally Isotropic Exceptional Projective

Homogeneous Varieties Are Locally Isotropic  491–500

Alena Pirutka and Nobuaki Yagita Note on the Counterexamples for the Integral Tate Conjecture

over Finite Fields  501–511

Vladimir L. Popov

Around the Abhyankar–Sathaye Conjecture  513–528

Anne Quéguiner-Mathieu and Jean-Pierre Tignol The Arason Invariant of Orthogonal Involutions of Degree 12 and 8,

and Quaternionic Subgroups of the Brauer Group  529–576

David J. Saltman Finite u Invariant and Bounds

on Cohomology Symbol Lengths  577–590